This is how it is written in my Math 11 textbook.
Baseball. The height, h(t) metres, of a batted baseball as a function of time, t seconds, since the ball was hit can be modelled by the function
h(t) = -2.1(t-2.4)^2 +13
what was its height when it was hit, to the nearest tenth of a metre?
What was the height of the ball, to the nearest tenth of a metre, 1 s after it was hit?








Answer by ewatrrr(24785)


Hi
h(t) metres represents the height of a batted baseball as a function of time, t seconds,
"since the ball was hit"
h(t) = -2.1(t-2.4)^2 + 13
h(0) = -2.1)-2.4)^2 + 13 = 1.1 to nearest tenth of metre
h(1) = -2.1(-1.4)^2 + 13 = 8.9 nearest tenth of metre